# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v1

@@ B.4.1 Uniform $(0,1)$
B.5 Some notes on the integral function $I(\alpha,\beta)$
In the derivation of the twCRPS for the log-normal distribution the integral function $I(\alpha,\beta)$ appears. It can be expressed in terms of Owen’s T function [Owen, 1956, 1980]:
$T(h,a)=\frac{1}{2\pi}\int_{0}^{a}\frac{e^{-\frac{1}{2}h^{2}(1+x^{2})}}{1+x^{2}},$ (53)
as:
@@ B.5 Some notes on the integral function $I(\alpha,\beta)$
$I(0,\beta)$ $=\Phi\left(\frac{\beta}{\sqrt{2}}\right)-\frac{1}{2}\Phi^{2}\left(\frac{\beta}{\sqrt{2}}\right),$ (56)
$I(0,0)$ $=\frac{3}{8}.$ (57)
There are routines available in R and other languages for evaluating Owen’s T function based on various series expansions or numerical integration. Alternatively we propose the following recursive power series expansion, using Hermite polynomials $H_{n}$:
@@
$J_{0}$ $=\frac{1}{2}[1-\Phi^{2}(\alpha)]$ (60)
@@
Using integration by parts and some properties of the Hermite polynomials yields the following recursive scheme. Initialize:
@@
$C_{n}$ $=-\frac{\alpha\beta}{n}C_{n-1}-\frac{n-3}{(n-1)n}\beta^{2}C_{n-2},\quad n\geq 3,$ (64)
$I_{n}$ $=C_{n}-\frac{n-2}{2(n-1)n}\beta^{2}I_{n}-2,\quad n\geq 3.$ (65)
In practice, the power series will be truncated after a finite number of terms as soon as $|I_{n}|$ has become small enough. By truncating the series after $N_{*}$ terms where $N_{*}$ is the smallest natural number such that $|I_{N_{*}-1}|<\eta$ and $|I_{N_{*}}|<\eta$ we have a rough estimate of the absolute error as:
$\text{est}=\max\left(\eta,\epsilon\max_{N=0,\dots,N_{*}}|\sum_{n=0}^{N}I_{n}|\right),$ (66)
where $\epsilon$ is the smallest number on the machine. In some numerical experiments we find that $I(\alpha,\beta)$ is evaluated rather fast and accurately using the above recursion formulas for any $\alpha$ and $\beta<7$. For other values the following recursive relationship can be derived using results from Owen [1980]:
$I(\alpha,\beta)=\Phi\left(\frac{\beta}{\sqrt{2}}\right)-\Phi(\alpha)+I\left(\frac{\beta}{\sqrt{2}},\sqrt{2}\alpha\right).$ (67)

### v2

@@ B.3.1 Continuous uniform distribution on $[a,b]$
B.3.2 Generalized Pareto distribution
twCRPS expression for $\tau\geq\mu$ and $\xi<1$:
$\text{twCRPS}_{\tau}\left(F_{\xi},y\right)=&v_{\tau}(y)-\tau+\frac{2}{1-\xi}\left\{{\left[1-F_{\xi}(v_{\tau}(y))\right]}^{1-\xi}-{\left[1-F_{\xi}(\tau)\right]}^{1-\xi}\right\}\\ &+\frac{1}{2-\xi}{\left[1-F_{\xi}(\tau)\right]}^{2-\xi},\qquad\qquad\tau\geq 0,\end{aligned}$
$\text{twCRPS}_{\tau}\left(F_{\xi,\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{\xi},\frac{y-\mu}{\sigma}\right),$ (46)
where $F_{\xi}$ is the CDF of the generalized Pareto distribution with shape parameter $\xi\in\mathbb{R}$, zero location and unit scale, and $F_{\xi,\mu,\sigma}(x)=F_{\xi}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
Source: derived by the authors.
B.3.3 Truncated logistic distribution on $[a,b]$
twCRPS expression for $a\leq\tau\leq b$ and $a\leq y\leq b$:
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=&\frac{1}{[F(b)-F(a)]^{2}}\left\{F(\tau)-F(b)+F^{2}(a)\log\frac{F(v_{\tau}(y))}{F(\tau)}\right.\\ &\left.+[1-F(a)]^{2}\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}+F^{2}(b)\log\frac{F(b)}{F(v_{\tau}(y))}+[1-F(b)]^{2}\log\frac{1-F(v_{\tau}(y))}{1-F(b)}\right\},\end{aligned}$
$\text{twCRPS}_{\tau}\left(F_{a,\mu,\sigma}^{b},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma},\frac{y-\mu}{\sigma}\right),$ (47)
where $F$ is the standard logistic CDF, $F_{a}^{b}$ is the standard logistic CDF truncated to $[a,b]$ with $-\infty<a<b<\infty$, and $F_{a,\mu,\sigma}^{b}(x)=F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
For only left truncation ($b\to\infty$), we have:
$\text{twCRPS}_{\tau}\left(F_{a}^{\infty},y\right)=&-\frac{1-F(\tau)}{[1-F(a)]^{2}}+\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}-\frac{F^{2}(a)}{[1-F(a)]^{2}}\log F(\tau)\\ &-\frac{1+F(a)}{1-F(a)}\log F(v_{\tau}(y)).\end{aligned}$
Source: For only left truncation, this formula is given by Allen et al. [2021b]; the general case of truncation to $[a,b]$ is derived by the authors.
B.3.4 Truncated normal distribution on $[a,b]$
twCRPS expression for $a\leq\tau\leq b$ and $a\leq y\leq b$:
$\begin{split}\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=&-\tau\left[\frac{\Phi(\tau)-\Phi(a)}{\Phi(b)-\Phi(a)}\right]^{2}+v_{\tau}(y)\left[2\frac{\Phi(v_{\tau}(y))-\Phi(a)}{\Phi(b)-\Phi(a)}-1\right]\\ &+\frac{2}{\Phi(b)-\Phi(a)}\left[\varphi(v_{\tau}(y))-\varphi(\tau)\frac{\Phi(\tau)-\Phi(a)}{\Phi(b)-\Phi(a)}\right]-\frac{1}{\sqrt{\pi}}\frac{\Phi(\sqrt{2}b)-\Phi(\sqrt{2}\tau)}{\left[\Phi(b)-\Phi(a)\right]^{2}},\end{split}$ (48)
$\text{twCRPS}_{\tau}\left(F_{a,\mu,\sigma}^{b},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma},\frac{y-\mu}{\sigma}\right),$ (49)
where $F_{a}^{b}$ is the standard normal CDF truncated to $[a,b]$ with $-\infty<a<b<\infty$, and $F_{a,\mu,\sigma}^{b}(x)=F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$. Here, $\varphi$ and $\Phi$ denote the standard normal PDF and CDF, respectively.
The case of only left truncation ($b\to\infty$) follows from $\Phi\left(\frac{b-\mu}{\sigma}\right)=\Phi\left(\frac{\sqrt{2}(b-\mu)}{\sigma}\right)=1$ for any $\mu\in\mathbb{R}$ and $\sigma>0$.
Source: This formula is given by Wessel et al. [2024].
B.4 Some notes on the integral function $I(\alpha,\beta)$
In the derivation of the twCRPS for the log-normal distribution the integral function $I(\alpha,\beta)$ appears (eq.(43)). It has the statistical interpretation as the probability $P(X>\alpha,Y<X+\beta)$ where $X$ and $Y$ are two independent standard normal random variables. The integral function $I$ can be expressed in terms of Owen’s $T$ function [Owen, 1956, 1980],
$T(h,a)=\frac{1}{2\pi}\int_{0}^{a}\frac{\exp\left[-\frac{1}{2}h^{2}(1+x^{2})\right]}{1+x^{2}},\qquad h\in\mathbb{R},\qquad a\in\mathbb{R},$ (50)
as
@@ B.4 Some notes on the integral function $I(\alpha,\beta)$
$I(0,\beta)$ $=\Phi\left(\frac{\beta}{\sqrt{2}}\right)-\frac{1}{2}\Phi^{2}\left(\frac{\beta}{\sqrt{2}}\right).$ (53)
There are routines available in R and other languages for evaluating Owen’s $T$ function based on various series expansions or numerical integration.
Alternatively, we propose a fast and stable method for evaluating $I(\alpha,\beta)$ based on a recursive power series expansion, making use of (probabilist’s) Hermite polynomials ${\{H_{n}\}}_{n=0}^{\infty}$. The integral function $I$ has the Taylor series
@@
with
$J_{0}=I(\alpha,0)=\frac{1}{2}[1-\Phi^{2}(\alpha)]$ (56)
and
@@
Using integration by parts and properties of the Hermite polynomials yields the following recursive scheme:
Initialize
@@
$C_{n}=$ $-\frac{\alpha\beta}{n}C_{n-1}-\frac{n-3}{(n-1)n}\beta^{2}C_{n-2},\qquad n\geq 3,$ (60)
$I_{n}=$ $C_{n}-\frac{n-2}{2(n-1)n}\beta^{2}I_{n-2},\qquad n\geq 3.$ (61)
In practice, the power series is truncated after a finite number of terms as soon as $|I_{n}|$ has become small enough. As the recursion may run in two branches of different magnitude involving even and odd indices $n$, respectively, we truncate the series after $N_{*}$ terms where $N_{*}$ is the smallest natural number such that $|I_{N_{*}-1}|<\eta$ and $|I_{N_{*}}|<\eta$ with, say, $\eta={10}^{-15}$. An estimate of the absolute error in $I(\alpha,\beta)$ which turns out to be very close to the actual absolute error can be obtained as
$\text{err}=\max\left(\eta,\varepsilon\max_{N=0,\dots,N_{*}}\left|\sum_{n=0}^{N}I_{n}\right|\right),$ (62)
where $\varepsilon$ is the smallest number on the machine, typically $\varepsilon=2^{-52}\approx 2.2\times{10}^{-16}$. In numerical experiments we find that $I(\alpha,\beta)$ is evaluated rather fast and accurately using the above recursion formulas for any $\alpha$ and about $|\beta|<7$. Computation can be further accelerated and the range of tractable values of $\beta$ extended by invoking the identity
$I(\alpha,\beta)=\Phi\left(\frac{\beta}{\sqrt{2}}\right)-\Phi(\alpha)+I\left(\frac{\beta}{\sqrt{2}},\sqrt{2}\alpha\right)$ (63)
which is derived using results from Owen [1980]. In summary, our scheme allows to evaluate the twCRPS for the log-normal distribution fast ($N_{\ast}<50$) and virtually to machine precision for any practically relevant combination of parameters $\mu$, $\sigma$ and $\tau$.